# Deep Generative Model for Human Mobility Behavior

### v1

@@ Abstract
Understanding and modeling human mobility is central to challenges in transport planning, sustainable urban design, and public health. Despite decades of effort, simulating individual mobility remains challenging because of its complex, context-dependent, and exploratory nature. Here, we present MobilityGen, a deep generative model that produces realistic mobility trajectories spanning days to weeks at large spatial scales. By linking behavioral attributes with environmental context, MobilityGen reproduces key patterns such as scaling laws for location visits, activity time allocation, and the coupled evolution of travel mode and destination choices. It reflects spatio-temporal variability and generates diverse, plausible, and novel mobility patterns consistent with the built environment. Beyond standard validation, MobilityGen yields insights not attainable with earlier models, including how access to urban space varies across travel modes and how co-presence dynamics shape social exposure and segregation. Our work establishes a new framework for mobility simulation, paving the way for fine-grained, data-driven studies of human behavior and its societal implications.
@@
Human mobility is inherently complex and often appears unstructured. From day to day, individuals visit locations across various spatial scales, using multiple transport modes at different periods of the day to engage in a range of activities. Indeed, movement decisions involve intricate psychological planning [Axhausen and Gärling, 1992] and are influenced by individual factors such as physical ability [Burbidge and Goulias, 2009] and socioeconomic status [Hanson and Hanson, 1981]. Despite this complexity, human mobility is fundamentally constrained by physical infrastructures [Jiang et al., 2009], such as road and rail networks, and heavily influenced by the surrounding environment [Handy et al., 2002, Hong et al., 2014]. Ample empirical evidence demonstrates that mobility patterns follow simple scaling laws [Brockmann et al., 2006, González et al., 2008, Alessandretti et al., 2018, Schläpfer et al., 2021], and exhibit striking regularities, for instance regarding spatial location choices [Song et al., 2010]. These collective regularities form the foundation for the realistic simulation of human mobility behaviors.
@@
Despite significant efforts and the importance of the task, simulating mobility behavior remains challenging. Current approaches to synthesizing disaggregated travel demand, such as the Equasim pipeline [Hörl and Balac, 2021], the OASIS framework [Pougala et al., 2023], and various other methods based on the principle of utility maximization [Bhat and Singh, 2000, Auld and Mohammadian, 2012], focus on the desires for activity and temporal constraints of individuals, while largely overlooking the spatial patterns of those activities. Conversely, state-of-the-art mechanistic models, including the Container model [Alessandretti et al., 2020], the exploration and preferential return (EPR) model [Song et al., 2010], and more recent variants derived from them [Pappalardo et al., 2015, Barbosa et al., 2015, Zhao et al., 2021], faithfully reproduce scaling laws of location visits, yet their limited modeling capacity restricts them from capturing the interactions between individuals’ spatial preferences and their other movement decisions [Pappalardo et al., 2023]. Neglecting those interactions can lead to misconceptions about individual space usage. This has far-reaching consequences for downstream applications that depend on accurate spatial modeling, such as sustainable transportation [Xu et al., 2018], human-centered urban design [Zhang et al., 2023], and epidemic prevention [Barreras and Watts, 2024].
@@
Data-driven deep learning models offer a possible solution to these challenges. In recent years, learning-based models have revolutionized image and text processing [LeCun et al., 2015]. Notably, natural language processing has experienced remarkable progress through new sequence processing techniques that excel at capturing complex patterns in large text collections, in particular, the transformer architecture [Vaswani et al., 2017] and the denoising diffusion probabilistic model (DDPM) [Sohl-Dickstein et al., 2015, Ho et al., 2020]. Their effectiveness in modeling language has facilitated applications across diverse domains concerned with spatial or temporal sequences, such as protein structures [Liu et al., 2024], turbulent flows [Li et al., 2024], or human life events [Savcisens et al., 2024]. Individual mobility can be conceptualized as a sequence of discrete, time-ordered events, each reflecting concurrent behavioral decisions made during movement. Still, language models are not directly applicable to human mobility behavior. Generalizing deep generative modeling techniques, originally designed for one-dimensional word sequences, to the multifaceted nature of mobility behavior shaped by external contexts, requires careful consideration.
@@
Here, we introduce MobilityGen, a deep generative model that simulates realistic, multifaceted human mobility over days and weeks and at large spatial scales. Recognizing that individual mobility unfolds as a sequence of events, MobilityGen leverages a DDPM to learn the generative mechanisms underlying these trajectories. Each event comprises multiple behavioral attributes, modeled as distinct but interdependent learning objectives, with their interactions captured through a shared latent representation that encodes meaningful structure across mode, time, and location. In addition, MobilityGen incorporates contextual features of the built environment. This allows the generation of realistic yet previously unseen behavioral variations. As our experiments show, the model reproduces national-scale visitation and transitions. It also captures behavioral dependencies that span extended temporal horizons, producing synthetic sequences closely aligned with real-world patterns. Together, these capabilities establish a unified framework for mobility simulation, enabling the generation of comprehensive, fine-grained mobility data at scale and supporting analyses previously not feasible with earlier models. This facilitates applications ranging from testing sustainable transport strategies and assessing accessibility in urban design to analyzing co-presence dynamics that influence social segregation and epidemic response.
Results
A deep generative model of individual mobility
@@ Results
To identify individual mobility patterns and establish a comprehensive model, we analyzed GNSS trajectories from a smartphone-based travel survey that recorded the mobility behavior of thousands of participants across Switzerland. We focused on $\approx$2100 individuals who were continuously tracked over an extended period with high spatio-temporal accuracy (see Methods). For subsequent sequence modeling, GNSS traces are segmented into stationary periods, i.e., fixed locations where individuals perform activities, and movement sections corresponding to transits between these activity locations. We then calculate activity attributes (such as time use and travel mode) and project the activity locations onto a hierarchical multi-resolution grid, with the highest resolution being $\approx$$\times500$m2 per cell (exact size varies slightly due to cartographic projection, see Methods). Inspired by activity-based modeling methods [Axhausen and Gärling, 1992], we conceptualize human mobility as sequences of activity events, each linked to specific attributes such as location, time use, and travel mode, along with contextual information (Figure 1a). We particularly focus on contexts that characterize the physical environment, including the coordinate geometry of locations and points of interest (POIs) that represent urban functional facilities.
@@
Figure 1: Modeling individual mobility with a deep generative framework. a. We conceptualize individual mobility as a sequence of chronologically ordered activity events, where each event has associated activity attributes (location, time use, and travel mode) and contextual information (coordinates and nearby POIs). b. Raw activity events are encoded into latent mobility embeddings through dedicated embedding modules. c. During training, the diffusion process adds Gaussian noise to the target sequence embedding $\mathbf{z}_{0}$. A transformer-based decoder learns to reverse this process and reconstruct a denoised embedding $\hat{\mathbf{z}}_{0}$. Features from the traveled sequence are extracted by an encoder and provided as guidance to the decoder. Final activity attributes are predicted via linear output heads. The model is trained by minimizing reconstruction and consistency losses between the generated and original sequences (see Methods). d. During inference, we sample a Gaussian noise vector $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively refine it using the decoder, guided by encoder-extracted features. The denoised embedding $\mathbf{z}_{0}$ is then decoded into the sequence of activity events. Travel mode icons adapted from Flaticon.com.
@@
The challenge for a comprehensive mobility model involves analyzing this multi-dimensional temporal data stream to derive realistic, multifaceted activity behavior. Traditional methods often focus on a single aspect, such as location choices [Alessandretti et al., 2020] or travel mode preferences [Reck et al., 2022], or are limited to describing simple behavioral dependencies [Jiang et al., 2016, Hörl and Balac, 2021]. In contrast, we propose MobilityGen, which jointly models these attributes and their relations with contextual information.
@@
MobilityGen employs learnable embedding modules to map raw attributes and contexts to a shared embedding space (Figure 1b). Learning comprehensive mobility embeddings enables MobilityGen to decouple and thoroughly explore the intricate structure of activity events. We utilize a DDPM [Sohl-Dickstein et al., 2015, Ho et al., 2020] to transform latent embeddings into event sequences. This recent class of generative models has emerged as a particularly powerful scheme to learn complex, high-dimensional data distributions, thus capable of modeling the mechanisms that govern these events. In a DDPM, the original sequence is gradually corrupted with tractable random noise, and an encoder-decoder network (in our case, a transformer [Vaswani et al., 2017]) learns the reverse denoising process (Figure 1c). Once trained, the (reverse) process can generate new, plausible sequences from pure Gaussian noise (Figure 1d). Additionally, MobilityGen uses a transformer encoder network to derive features from the observed traveled sequence, acting as a guiding signal during the denoising process for generating event embeddings. Finally, linear decoder heads extract the respective activity attributes from the predicted embeddings. Together, these attributes form the final output, namely a new, simulated sequence of activity events for an individual, sampled in a way that respects the observed transition and combination patterns of activity attributes.
Microscopic modeling reproduces location choice behavior
To demonstrate that the newly generated activity events replicate realistic mobility behaviors, we begin by evaluating the ability of MobilityGen to capture location patterns. How individuals select spatial locations for different activities shapes their interactions with physical space and is a fundamental driver of many urbanization-induced challenges [Moro et al., 2021, Nilforoshan et al., 2023, Liu et al., 2023]. Here, for every traveled sequence in the test dataset, we allow the trained MobilityGen to generate a synthetic sequence containing 50 activity events (see Methods). We focus on the simulated visits to locations and compare them with the real visits from empirically observed sequences.
We first assess the ability of MobilityGen to capture fundamental properties of location choices and compare its performance against state-of-the-art microscopic mobility models. This is done through representative statistical location metrics, including the distribution of location visit frequency, the evolution of the radius of gyration, and the distribution of mobility entropy (see Methods). We evaluate each metric for real data and for location traces generated by MobilityGen, EPR [Song et al., 2010], and Container [Alessandretti et al., 2020] (Figure 2). We begin with the average rank distribution of location visits (Figure 2a), a widely used measure of preference for activity locations [González et al., 2008, Barbosa et al., 2018]. Our results indicate that MobilityGen provides a significantly better fit to real data than both the EPR and Container models, as confirmed by the likelihood ratio test [Clauset et al., 2009] (with $P<10^{-3}$) and the Wasserstein distances [Panaretos and Zemel, 2019] between the rank distributions (see Extended Data Table 1). The radius of gyration, which measures the characteristic distance an individual covers [González et al., 2008], serves to quantify the movement intensity in space. We observe that the median evolution of the radius of gyration across individuals, $r_{u}(k)$, over displacement steps $k$, is well described by logarithmic growth for both real and generated location sequences (Figure 2b), consistent with previous findings [Alessandretti et al., 2020]. However, when fitting $r_{u}(k)\propto\alpha\log(k)$, only the value $\alpha$ derived from MobilityGen agrees well with the real data (see Supplementary Note 1). Finally, we evaluate the temporal mobility entropy that quantifies the predictability of location sequences, considering both the frequency and the order of visits [Song et al., 2010]. Using the likelihood ratio test and the Wasserstein metric, we further confirm that, compared against the EPR and Container models, MobilityGen most accurately approximates the true entropy distribution over the population (with $P<10^{-3}$, see Figure 2c and Extended Data Table 1). Additionally, we evaluated a representative set of state-of-the-art data-driven mobility models and further location metrics (see Supplementary Note 1.1). Our findings indicate that MobilityGen consistently provides the best fit to real-world data. Further evaluations with metrics from language modeling confirm that MobilityGen can simulate location visits that closely resemble the real sequence while avoiding regurgitation of input sequences and maintaining sufficient diversity (see Supplementary Note 1.2).
To further evaluate the generated mobility patterns, we examine the spatial distribution of location visits. As an illustrative example, we map the observed and synthetic trajectories of a sample individual (Figure 3a and b). The comparison shows that MobilityGen effectively captures both visit and transition patterns at frequently visited locations (thick lines), while also generating occasional visits to less familiar locations (thin lines). Accurately modeling such exploratory behavior remains a central challenge in human mobility research [Cuttone et al., 2018]. To assess how well MobilityGen addresses this challenge at the population level, we compare the aggregated spatial distribution of location visits across Switzerland in the real and simulated data (Figure 3c and d). Visit frequencies are color-coded from low (yellow) to high (red). MobilityGen closely matches the concentration of frequently visited locations and also generates visits to less frequented areas, reflecting a balance between individual-level exploration and population-level spatial preference. Using the entropy $E=-\sum_{k}p(f_{k})\log p(f_{k})$ to quantify the dispersion of location visits $f_{k}$, we obtain 0.33 for the real data and 0.53 for the simulated data, indicating greater spatial dispersion in the latter. To analyze spatial dynamics at a finer resolution, we examine three representative cities: Bern, Zurich, and Lucerne (Figure 3e-j), each characterized by distinct urban functions and mobility patterns. MobilityGen successfully reproduces key urban visitation structures in all three cases. However, spatial dispersion remains higher in the simulated data, with entropy values of 2.49 vs. 1.89 for Bern, 3.99 vs. 3.39 for Zurich, and 2.71 vs. 2.11 for Lucerne. Despite this discrepancy, MobilityGen compares favorably to other microscopic mobility models in replicating the location visit distributions (Extended Data Table 2). Complementary visual comparisons are provided at both national and urban levels in Extended Data Figure 1 and 2.
Structured behavioral patterns emerge across time, mode, and location
To move beyond location accuracy, we frame the mobility generation task as a multi-task learning problem that enables MobilityGen to jointly model a diverse set of behavioral attributes, including activity timing, travel mode, and interdependencies among attributes (see Methods). We evaluate the model’s ability to capture these structured patterns by generating 50 activity events for each sequence in the test set, focusing on behavioral regularities and attribute interactions (Figure 4).
Incorporating temporal information allows MobilityGen to reason about the scheduling of each activity. This temporal dimension is evident in the activity durations (Figure 4a) and the number of daily visited locations (Figure 4b), both of which closely match those observed in the real data. To further examine patterns of activity participation, we construct mobility motifs [Schneider et al., 2013], which serve as high-level descriptors of regularities in activity location and scheduling. These motifs refer to unique topological structures that emerge when daily location visits are represented as graphs, with locations as nodes and trips between them as directed edges. From the real sequences, we identify eight motifs that occur on more than 0.5% of days (see Methods). MobilityGen accurately recovers the relative frequencies of both common and rare motifs, including simple two-location transitions and more complex patterns involving up to five locations (Figure 4c). We further benchmark MobilityGen against two state-of-the-art spatio-temporal mobility simulators, the Diary-based Trajectory Simulator (DITRAS) [Pappalardo and Simini, 2018] and TimeGeo [Jiang et al., 2016], using additional mobility metrics (see Extended Data Table 3 and Supplementary Note 2.1). Across all measures, MobilityGen more faithfully reproduces the interplay between activity timing and location choice, yielding the closest alignment with empirical behavior.
Next, we incorporate user-verified travel mode labels into MobilityGen to jointly model mode choice in relation to activity events. Figure 4d compares the travel mode distributions from the observed and simulated activity sequences, with discrepancies quantified by the Wasserstein distance (see Extended Data Table 3). Across all scenarios, MobilityGen successfully reconstructs the mode preferences exhibited in real-world behavior. Car and walking, the two most common modes, show excellent agreement between empirical and simulated shares. Notably, MobilityGen also captures subtler distinctions among public modes, including bus, train, and tram, accurately reflecting their relative frequencies. We further assess mode-specific behavior by comparing the displacement distributions for car and walking trips (Figure 4e, f). For car travel, the simulated displacements closely align with the observed distribution (Kolmogorov–Smirnov statistic $KS=0.02$, with $P<10^{-3}$). For walking, MobilityGen produces slightly longer displacements ($KS=0.06$, $P<10^{-3}$), potentially due to short trips being disproportionately affected by spatial discretization. This effect could be mitigated in future work by adopting a finer location grid, although such refinement may increase computational complexity.
Finally, we assess the evolution of trip packages, defined as combinations of travel modes and activity locations [Hong et al., 2023], to evaluate MobilityGen’s ability to model complex attribute interactions (Figure 4g; see Methods). The average number of unique trip packages over time, denoted as $Tp(t)$, follows a sub-linear growth over time $t$, consistent with previous findings [Hong et al., 2023]. The scaling exponent $\alpha$ in the relation $Tp(t)\propto t^{\alpha}$ is nearly identical between simulated ($\alpha=0.85\pm 0.01$) and empirical sequences ($\alpha=0.85\pm 0.01$), indicating that MobilityGen accurately captures the evolving diversity of travel mode and activity location choices over time. Further ablation studies confirm that the full multi-attribute model achieves a more realistic representation of attribute interactions than simplified variants with fewer modeled attributes (see Supplementary Note 3.1).
To understand how MobilityGen captures attribute interactions and generates realistic activity sequences, we analyze the structure of its learned embedding space. We apply densMAP [Narayan et al., 2021] to project the 128-dimensional embeddings into two dimensions for visualization (see Supplementary Note 2.2). The projections reveal that MobilityGen learns semantically meaningful representations: similar travel modes appear near each other, temporal attributes vary smoothly along continuous trajectories, and frequently visited locations are well-separated from infrequent ones. When multiple attributes are jointly embedded, the space reveals distinct event clusters tied to specific locations, alongside a more diffuse region that likely reflects less location-dependent activities. These patterns suggest that complex mobility dynamics can be inferred directly from behavioral sequences, with the embedding space offering a compact and interpretable representation that supports diverse mobility analysis.
Context enhances generalization to novel locations
A major feature of MobilityGen is its capacity to let contextual information influence individual travel decisions – an aspect well-supported by the literature but not adequately covered by existing microscopic mobility models [Pappalardo et al., 2023]. In MobilityGen, these contexts are flexibly combined with activity attributes through learnable parameters, resulting in an implicit context-aware mobility model (Figure 1b, and see Methods). We restrict our analysis to spatial contexts that characterize the physical environment, namely the coordinate geometry, which reveals movement directionality [Zhao et al., 2024]; and POIs, which display urban functions and imply movement purposes [Hong et al., 2023]. Further contexts could be integrated through appropriate representations; we defer this to future research.
We analyze location patterns to assess the influence of contextual factors on the simulation performance of MobilityGen. Generated locations are compared to observed data and categorized into three sets: Set A includes locations present in both sequences, while Sets B and C contain locations visited exclusively in the observed and generated sequences (Figure 5a). Since MobilityGen replicates key location preferences (see Figure 3), Set A is expected to capture the locations where individuals conduct most of their activities. Indeed, an analysis of the visit frequencies reveals that 96.4% of the true visits and 93.4% of the generated visits occur at locations in Set A (Figure 5b). In contrast, locations in Sets B and C receive relatively few visits, reflecting individuals’ tendencies toward novelty-seeking and the exploration of unfamiliar locations [Arentze and Timmermans, 2005, Song et al., 2010, Alessandretti et al., 2018].
To understand how MobilityGen models novelty-seeking behavior, we identify the functional types of locations using POI data. Applying $k$-means clustering to POI representations of locations yields five distinct location classes, whose functional types are inferred from their spatial distributions and the enrichment of specific POI categories [Yao et al., 2017, Niu and Silva, 2021] (see Methods). This functional classification enables a detailed analysis of location preferences, revealing that visits in Set A predominantly occur in the City center, whereas visits in Sets B and C are largely concentrated in Residential & Countryside. Moreover, preferences for Leisure & Recreation and Nature & Tourism sites are markedly higher in Sets B and C than in Set A. Using pairwise Wasserstein distances to quantify the similarity of visitation patterns across sets, we find that Sets B and C are closely aligned (Wasserstein distance: 0.11), while their distributions differ substantially from that of Set A (Set A vs. Set B: 0.57; Set A vs. Set C: 0.56). In other words, MobilityGen effectively captures patterns of exploratory behavior, despite the inherent difficulty of characterizing novel locations and predicting which specific places individuals are likely to explore.
A comparison with a model that excludes contextual information confirms that incorporating context improves the ability to characterize locations (see Supplementary Note 3.1). This improvement is particularly relevant in settings where historical mobility information is limited, such as when individuals seek novel locations for their activities. To benchmark how effectively MobilityGen leverages contextual cues, we conducted comparative experiments against established mobility flow generation models: Gravity [Zipf, 1946], Deep Gravity [Simini et al., 2021], and random forest [Cabanas-Tirapu et al., 2025]. These baselines estimate flows between location pairs using features such as distance, population, and selected facility densities (e.g., shops and transport hubs). Flow-based models differ fundamentally from MobilityGen in both problem formulation and modeling capacity. For instance, they cannot capture temporal dynamics or disaggregated mobility. To ensure a fair comparison, we focused exclusively on locations that were not observed during training (see Supplementary Note 3.2). The results show that MobilityGen outperforms the flow generation baselines by a substantial margin across multiple flow metrics (see Extended Data Table 4). We attribute this improved performance in novel locations to the integration of spatial features, including coordinate geometry and detailed POI information, as well as the joint modeling of contextual and behavioral factors. While further research is needed to fully understand novelty-seeking mobility, our findings support the hypothesis that comprehensive context modeling is necessary to capture the variability in human behavior.
Reproducing mode-specific spatial behavior and co-presence dynamics
Beyond standard validation, we ask whether the model enables analyses that have so far been difficult or infeasible with existing simulators. Our simulations generate realistic daily schedules that jointly capture where people go, how they travel, and when activities occur. Earlier tools often rely on predefined rules or cover only part of this behavioral structure, which has limited their use for studying differences across travel modes or patterns of shared presence in everyday settings. We focus on two representative scenarios with broad relevance to transport planning and urban research: (i) mode-specific spatial visitation, showing how individuals use different travel modes to move through urban space, and (ii) spatio-temporal co-presence, reflecting how individuals co-occur in space and time within shared environments.
While spatial visitation captures where individuals go, co-presence patterns reveal whom they encounter, offering complementary insight into the intensity and distribution of interpersonal contact in daily urban life. A key application of co-presence lies in socio-spatial segregation research, where attention has shifted from static residential indicators to individual-level experienced segregation enabled by fine-grained mobility data [Liao et al., 2025]. We assess MobilityGen’s ability to reproduce these patterns by focusing on income-based segregation. Following Moro et al. [2021], we group individuals into four income quantiles and measure, at each location, the time spent by people from different income groups (see Methods). We then calculate, for each individual, a segregation score that reflects their exposure to others across income levels. Simulated and observed segregation scores show strong agreement (Pearson $r=0.70$, $P<10^{-3}$; Figure 7), capturing both overall trends and the variability in co-presence-based income segregation. A breakdown by income quantiles confirms that MobilityGen reproduces systematic differences in exposure and substantially outperforms baseline simulators such as TimeGeo and DITRAS, whose separation of space and time limits their ability to capture realistic contact patterns (see Extended Data Figure 4). This case study illustrates MobilityGen’s broader value as a generative model of social exposure. By capturing individual-level interactions in daily schedules, it provides a simulation environment to explore how population-level social mixing evolves under different scenarios. This extends beyond segregation analysis to domains such as information diffusion and infectious disease spread, where spatial co-presence fundamentally shapes outcomes.
Discussion
Despite growing interest in understanding human movement patterns and their implications, existing simulation models fail to capture the full spatio-temporal variability of mobility behavior. Here, we address this gap with MobilityGen, a deep generative model that produces realistic sequences of mobility events with detailed behavioral attributes. MobilityGen consistently outperforms baseline models across key mobility indicators and, crucially, generates novel behavioral interactions across time, space, and mode that earlier approaches cannot capture. It also enables analyses that were previously infeasible. By linking behavioral attributes with contextual information, MobilityGen uncovers how urban space is accessed differently across travel modes and how co-presence dynamics shape patterns of social exposure, such as income-based segregation. These examples show how generative modeling moves beyond reproducing standard metrics to open new avenues for research and policy enabled by fine-grained mobility data.
Our data-driven framework introduces a new paradigm for modeling individual mobility behavior. By integrating behavioral attributes into a shared latent space, MobilityGen captures their interactions across extended temporal and spatial scales. This facilitates the discovery of complex, longitudinal, and nuanced behavioral patterns that are otherwise difficult to model using traditional approaches. Looking ahead, MobilityGen could be combined with utility-based activity models in a hybrid approach, improving interpretability and allowing simulations that account for household activity participation and targeted mobility interventions. Its flexibility also supports the integration of richer contextual information, from detailed socio-demographic characteristics to social networks and environmental factors such as weather. Together, these extensions position MobilityGen as a bridge between mobility research and disciplines such as social, health, and environmental sciences, representing a step toward an integrative framework for interdisciplinary mobility modeling.
Methods
Task formulation: activity event sequence generation

### v2

@@ Abstract
Understanding and modeling human mobility is central to challenges in transport planning, sustainable urban design, and public health. Despite decades of effort, simulating individual mobility remains challenging because of its complex, context-dependent, and exploratory nature. Here, we advance a unified event-level formulation of daily mobility and propose MobilityGen to generate multi-attribute event sequences over days to weeks at large spatial scales. By linking behavioral attributes with environmental context, MobilityGen reproduces key patterns such as scaling laws for location visits, activity time allocation, and the coupled evolution of travel mode and destination choices. It reflects spatio-temporal variability and generates diverse and plausible mobility patterns consistent with the built environment. Beyond standard validation, MobilityGen enables analyses that have been difficult with earlier models, including how access to urban space varies across travel modes and how co-presence dynamics shape social exposure and segregation. Together, these results support an integrated, data-driven basis for fine-grained studies of human mobility behavior and its societal implications.
keywords:
human mobility; activity schedules; deep generative model; transport planning; travel behavior modeling
@@
Human mobility is inherently complex and often appears unstructured. From day to day, individuals visit locations across various spatial scales, using multiple transport modes at different periods of the day to engage in a range of activities. Indeed, movement decisions involve intricate psychological planning [Axhausen and Gärling, 1992] and are influenced by individual factors such as physical ability [Burbidge and Goulias, 2009] and socioeconomic status [Hanson and Hanson, 1981]. Despite this complexity, human mobility is fundamentally constrained by physical infrastructures [Jiang et al., 2009], such as road and rail networks, and heavily influenced by the surrounding environment [Handy et al., 2002, Hong et al., 2014]. Ample empirical evidence demonstrates that mobility patterns follow simple scaling laws [Brockmann et al., 2006, González et al., 2008, Alessandretti et al., 2018, Schläpfer et al., 2021], and exhibit striking regularities, for instance regarding spatial location choices [Song et al., 2010a]. These collective regularities form the foundation for the realistic simulation of human mobility behaviors. For transport planning, the key challenge is to simulate not only where people travel, but also when activities occur, and how trips are made, while preserving their interdependencies across days.
@@
Despite significant efforts and the importance of the task, simulating mobility behavior remains challenging. Current approaches to synthesizing disaggregated travel demand, such as the Equasim pipeline [Hörl and Balac, 2021], the OASIS framework [Pougala et al., 2023], and various other methods based on the principle of utility maximization [Bhat and Singh, 2000, Auld and Mohammadian, 2012], focus on activity participation and temporal constraints, while often simplifying the spatial structure of activity locations and travel. Conversely, state-of-the-art mechanistic models, including the Container model [Alessandretti et al., 2020], the exploration and preferential return (EPR) model [Song et al., 2010b], and more recent variants derived from them [Pappalardo et al., 2015, Barbosa et al., 2015, Zhao et al., 2021], faithfully reproduce scaling laws of location visits, yet their limited modeling capacity restricts them from capturing the interactions between individuals’ spatial preferences and their other movement decisions [Pappalardo et al., 2023]. Neglecting those interactions can lead to misconceptions about individual space usage. This has far-reaching consequences for downstream applications that depend on accurate spatial modeling, such as sustainable transportation [Xu et al., 2018], human-centered urban design [Zhang et al., 2023], and epidemic prevention [Barreras and Watts, 2024].
@@
Data-driven generative models offer a possible route to simulation that preserves multi-attribute travel behavior at scale. In recent years, learning-based models have revolutionized image and text processing. Notably, natural language processing has experienced remarkable progress through new sequence processing techniques that excel at capturing complex patterns in large text collections, in particular, the transformer architecture [Vaswani et al., 2017] and the denoising diffusion probabilistic model (DDPM) [Sohl-Dickstein et al., 2015, Ho et al., 2020]. Their effectiveness in modeling language has facilitated applications across diverse domains concerned with spatial or temporal sequences, such as protein structures [Liu et al., 2024], turbulent flows [Li et al., 2024], or human life events [Savcisens et al., 2024]. Inspired by these advances, recent work has begun to adapt deep generative frameworks to human mobility and trajectory synthesis [Chu et al., 2024, Song et al., 2024, Long et al., 2025]. These studies highlight the promise of data-driven generation, while leaving open the challenge of jointly modeling the interdependent spatio-temporal and modal structure of daily mobility behavior at scale. Individual mobility is naturally conceptualized as a sequence of discrete, time-ordered events, each reflecting concurrent behavioral decisions made during movement. Still, language models are not directly applicable to human mobility behavior. Generalizing deep generative modeling techniques, designed for one-dimensional word sequences, to multifaceted, context-dependent mobility behavior requires careful consideration.
@@
Here, we introduce MobilityGen, which advances a unified generative view of daily mobility as sequences of multi-attribute events unfolding over days and weeks at large spatial scales. In this view, where people go, when activities occur, and how they travel are modeled as coupled behavioral decisions rather than separable components. We instantiate this formulation with a DDPM to learn the structure of multi-attribute event sequences from data. Each event comprises multiple behavioral attributes, modeled as distinct but interdependent learning objectives, with their interactions captured through a shared latent representation that encodes meaningful structure across mode, time, and location. In addition, MobilityGen incorporates contextual features of the built environment, enabling the generation of diverse behavioral variations consistent with observed constraints and context. As our experiments show, MobilityGen preserves national-scale visitation and transition regularities. It also captures behavioral dependencies that span extended temporal horizons, producing synthetic sequences aligned with real-world patterns. Together, these results demonstrate the utility of this event-level formulation for simulating coupled daily behavior at scale. This enables the generation of comprehensive, fine-grained mobility data and supports applications ranging from testing sustainable transport strategies and assessing accessibility in urban design to analyzing co-presence dynamics that influence social segregation and epidemic response.
1 Methodology
1.1 A coupled event-level representation of individual mobility
@@
To characterize individual mobility and support a comprehensive event-level model, we analyzed global navigation satellite system (GNSS) trajectories from a smartphone-based travel survey that recorded the mobility behavior of thousands of participants across Switzerland. GNSS traces are segmented into stationary periods, i.e., fixed locations where individuals perform activities, and movement sections corresponding to transits between these activity locations. We then calculate activity attributes (such as time use and travel mode) and project the activity locations onto a hierarchical multi-resolution grid, with the highest resolution being $\approx$$\times500$m2 per cell. Inspired by activity-based modeling methods [Axhausen and Gärling, 1992], we conceptualize human mobility as sequences of activity events, each linked to specific attributes such as location, time use, and travel mode, along with contextual information (Figure 1a). We particularly focus on contexts that characterize the built environment, including the coordinate geometry of locations and points of interest (POIs) that represent urban functional facilities.
@@
Figure 1: Modeling individual mobility with an event-level generative framework. a. We conceptualize individual mobility as a sequence of chronologically ordered activity events, where each event has associated activity attributes (location, time use, and travel mode) and contextual information (coordinates and nearby POIs). b. Raw activity events are encoded into latent mobility embeddings through dedicated embedding modules. c. During training, the diffusion process adds Gaussian noise to the target sequence embedding $\mathbf{z}_{0}$. A transformer-based decoder learns to reverse this process and reconstruct a denoised embedding $\hat{\mathbf{z}}_{0}$. Features from the traveled sequence are extracted by an encoder and provided as guidance to the decoder. Final activity attributes are predicted via linear output heads. The model is trained by minimizing reconstruction and consistency losses between the generated and original sequences (see Methods). d. During inference, we sample a Gaussian noise vector $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively refine it using the decoder, guided by encoder-extracted features. The denoised embedding $\mathbf{z}_{0}$ is then decoded into the sequence of activity events. Travel mode icons adapted from Flaticon.com.
@@
The challenge for a comprehensive mobility model requires analyzing this multi-dimensional temporal data stream to derive realistic, multifaceted activity behavior. Traditional methods often focus on a single aspect, such as location choices [Alessandretti et al., 2020] or travel mode preferences [Reck et al., 2022], or are limited to describing simple behavioral dependencies [Jiang et al., 2016, Hörl and Balac, 2021]. In contrast, MobilityGen advances a unified event-level formulation that jointly models these attributes and their interactions with contextual information.
@@
MobilityGen employs learnable embedding modules to map raw attributes and contexts to a shared embedding space (Figure 1b). Learning comprehensive mobility embeddings enables the model to encode the rich multi-attribute dependencies of activity events. We utilize a DDPM [Sohl-Dickstein et al., 2015, Ho et al., 2020], a powerful scheme for learning complex, high-dimensional distributions, to capture the structure of mobility event sequences. In a DDPM, the original sequence is gradually corrupted with tractable random noise, and an encoder-decoder network (in our case, a transformer [Vaswani et al., 2017]) learns the reverse denoising process (Figure 1c). Once trained, the (reverse) process can generate new, plausible sequences from pure Gaussian noise (Figure 1d). Additionally, a transformer encoder derives features from the observed travel sequence, acting as a guiding signal during denoising for generating event embeddings. Finally, linear decoder heads extract the respective activity attributes from the predicted embeddings. Together, these attributes form the final output, namely a new, simulated sequence of activity events for an individual, sampled in a way that respects observed transition and combination patterns across attributes.
1.2 Problem definition